# A New Algorithm for Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons

### v1

@@ 3.3 Convergence Analysis of Iterative Sequence
Before presenting the next theorem, we first introduce two families of linear operators, $\mathcal{L}^{*}_{P^{(k)}}$ and $\mathcal{L}^{(k,k+1)}$.
$\mathcal{L}^{*}_{P^{(k)}}(Y)=\,Y\left(A_{(0)}+B_{1}\hat{K}_{1}(P^{(k)})+B_{2}L)\right)+\left(A_{(0)}+B_{1}\hat{K}_{1}(P^{(k)})+B_{2}L)\right)^{\top}Y$ (29)
$+\sum_{l=1}^{r}\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P^{(k)})+D_{l,2}L)\right)^{\top}Y\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P^{(k)})+D_{l,2}L\right)\,,\quad\forall\,Y\in\mathbb{S}^{n},k=0,1,2,\ldots$
$\mathcal{L}^{(k,k+1)}(Y)=\,Y\left(A_{(k)}+B_{2}T_{(k)}\right)+\left(A_{(k)}+B_{2}T_{(k)}\right)^{\top}Y$ (30)
$+\sum_{l=1}^{r}\left(C_{l,(k)}+D_{l,2}T_{(k)}\right)^{\top}Y\left(C_{l,(k)}+D_{l,2}T_{(k)}\right)\,,\quad\forall\,Y\in\mathbb{S}^{n},k=0,1,2,\ldots$
where $T_{(k)}=-R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}$, $A_{(k)},C_{l,(k)}(1\leq l\leq r),N_{2,(k)}$ is defined in (11) and (12).
@@
  • There exist matrices $E_{l,(0)}(1\leq l\leq r)$ such that: $\sum_{l=0}^{r}E_{l,(0)}^{\top}E_{l,(0)}=Q-S^{\top}(0)R(0)^{-1}S(0)$ and the system
    $\left[E_{0,(0)},E_{1,(0)},\dots,E_{r,(0)};A_{(0)},C_{1,(0)},\ldots,C_{r,(0)}\right]$
@@
The sequences $\{Z^{(k)}\}_{k\geq 0}$, $\{P^{(k)}\}_{k\geq 0}$ are well defined by (8) (9) (10) , and for each $k=0,1,2\ldots$ the following items are fulfilled:
@@
    $\text{Spec}\,\mathcal{L}^{(k,k+1)}\subset\mathbb{C}^{-}$;
@@
    Let $L\in\mathcal{A}$ is arbitrary but fixed and $\tilde{P}_{L}$ be the stabilizing solution to the corresponding ARE (7), then $\tilde{P}_{L}\succeq P^{(k)}+Z^{(k)}$ and $\text{Spec}\,\mathcal{L}_{P^{(k+1)}}\subset\mathbb{C}^{-}$;
@@
    $P^{(k)},P^{(k)}+Z^{(k)}\in\text{Dom }\mathcal{G}$ and $\mathcal{G}(P^{(k)}+Z^{(k)})=-\left(N_{1,(k)}-R_{12}(P^{(k)}+Z^{(k)})R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}\right)^{\top}$
    $\times\mathbb{R}_{22}^{\sharp}(P^{(k)}+Z^{(k)})^{-1}\left(N_{1,(k)}-R_{12}(P^{(k)}+Z^{(k)})R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}\right),$
    where $\mathbb{R}_{22}^{\sharp}(P^{(k)}+Z^{(k)}),N_{1,(k)},N_{2,(k)}$ is defined in (12);
$\lim_{k\rightarrow\infty}P^{(k)}=P^{*}$, where $P^{*}$ is the stabilizing solution of the stochastic GTARE (3).
@@
When $k=0$ , $P^{(0)}=0$ and $Z^{(0)}$ is the stabilizing solution to the following ARE:
$ZA_{(0)}+A_{(0)}^{\top}Z+\sum_{l=1}^{r}C_{l,(0)}^{\top}ZC_{l,(0)}+Q-S^{\top}(0)R(0)^{-1}S(0)$ (31)
$-\left(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(0)}\right)^{\top}\left(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,2}\right)^{-1}\left(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(0)}\right)=0.$
To show that $Z^{(0)}$ is well-defined, we first prove the existence unique stabilizing solution to the above equation. Since the set $\mathcal{A}$ is nonempty and set $L^{(0)}=L$, we can find the system $(A_{(0)}+B_{2}L^{(0)},C_{1,(0)}+D_{1,2}L^{(0)},\dots,C_{r,(0)}+D_{r,2}L^{(0)})$ associated with the system $(A_{L},C_{1L},\dots,C_{rL})$ is mean-square stable. This means the system
@@
is mean-square stabilizable. Since the system $E_{0,(0)},E_{1,(0)},\dots,E_{r,(0)};A_{(0)},C_{1,(0)},\dotsb,C_{r,(0)}]$ is stochastically detectable. And the assumption $R_{22}\succ 0$ and $Q-S^{\top}(0)R(0)^{-1}S(0)\succ 0$ ensures that
@@
Consequently, by Proposition 3.7, the Eq. (31) admits a unique stabilizing solution $Z^{(0)}$ such that $R_{22}(Z^{(0)})\succ 0$. Thus, $Z^{(0)}$ is well-defined as the unique stabilizing solution to the equation. This further implies that $\text{Spec}\,\mathcal{L}^{(0,1)}\subset\mathbb{C}^{-}$.
Let $L\in\mathcal{A}$ be arbitrary but fixed, $\tilde{P}_{L}$ denote the stabilizing solution to the corresponding ARE (7). Since the operator $\mathcal{L}^{*}_{P^{(0)}}$ is associated with the system $(A_{L},C_{1L}\dotsc,C_{rL})$ is mean-square stable, we have the $\text{Spec}\,\mathcal{L}_{P^{(0)}}\subset\mathbb{C}^{-}$. From (i.) and (ii.) in Lemma 3.8, we get $\tilde{P}_{L}\succeq P^{(0)}+Z^{(0)}$ and $\text{Spec}\,\mathcal{L}_{P^{(1)}}\subset\mathbb{C}^{-}$. Combining $R_{22}(\tilde{P}_{L})\succeq R_{22}\succ 0$ and $R_{11}\preceq R_{11}(\tilde{P}_{L})\prec 0$, we have $P^{(0)}$ and $\tilde{P}_{L}\in\text{Dom }\mathcal{G}$. Since $R_{22}(P^{(0)}+Z^{(0)})\succeq R_{22}\succ 0$ and $R(P^{(0)}+Z^{(0)})\prec R(\tilde{P}_{L})$, we obtain that $R_{11}(P^{(0)}+Z^{(0)})\preceq\mathbb{R}_{22}^{\sharp}(P^{(0)}+Z^{(0)})\preceq\mathbb{R}_{22}^{\sharp}(\tilde{P}_{L})\prec 0$ by using Corollary 4.5 in [16]. It follows that $P^{(0)}+Z^{(0)}\in\text{Dom }\mathcal{G}$. Substituting Eq. (8) into $\mathcal{G}\left(P^{(0)}+Z^{(0)}\right)$ by using Proposition 3.3, then gives:
$\mathcal{G}\left(P^{(0)}+Z^{(0)}\right)=(N_{1,(0)}-R_{12}(Z^{(0)})R_{22}(Z^{(0)})^{-1}N_{2,(0)})^{\top}R^{\sharp}_{22}(Z^{(0)})^{-1}(N_{1,(0)}-R_{12}(Z^{(0)})R_{22}(Z^{(0)})^{-1}N_{2,(0)}).$
@@
Assume k=h1k=h-1, $Z^{(h-1)}$ is well-defined and that $a_{h-1}-c_{h-1}$ hold. We shall prove that $Z^{(h)}$ is well-defined and that $a_{h}-c_{h}$ hold.
@@
$ZA_{(h)}+A_{(h)}^{\top}Z+\sum_{l=1}^{r}C_{l,(h)}^{\top}ZC_{l,(h)}+\sum_{l=0}^{r}E^{\top}_{l,(h)}E_{l,(h)}$ (32)
$-\left(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(h)}\right)^{\top}\left(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,2}\right)^{-1}\left(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(h)}\right)=0,$
where
$E_{l,(h)}=\begin{bmatrix}\mathbb{O}_{m_{1}\times n}^{\times(l)}\\ \sqrt{-\frac{1}{r+1}\mathbb{R}_{22}^{\sharp}\left(P^{(h)}\right)^{-1}}\left(N_{1,(h-1)}-R_{12}(P^{(h)})R_{22}(P^{(h)})^{-1}N_{2,(h-1)}\right)\\ \mathbb{O}_{m_{1}\times n}^{\times(r-l)}\end{bmatrix}\in\mathbb{R}^{[m_{1}(r+1)]\times n},0\leq l\leq r.$
We demonstrate that $Z^{(h)}$ is well-defined. By the Proposition 3.7, proving that $Z^{(h)}$ is well-defined is equivalent to proving that the system $\left[A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)};B_{2},D_{1,2},\dotsb,D_{r,2}\right]$ is mean-square stabilizable and $\left[E_{0,(h)},E_{1,(h)},\dots,E_{r,(h)};A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)}\right]$ is stochastically detectable. To prove the stability of the system $A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)};B_{2},D_{1,2},\dotsb,D_{r,2}]$, it suffices to set $L^{(h)}=L-\hat{K}_{2}(P^{(h)})$, where $\hat{K}_{2}(P)$ is defined in (21). We can then identify the system $\left(A_{(h)}+B_{2}L^{(h)},C_{1,(h)}+D_{1,2}L^{(h)},\dots,C_{r,(h)}+D_{r,2}L^{(h)}\right)$ associated with the operator $\mathcal{L}_{P^{(h)}}$. Since the condition $\text{Spec}\,\mathcal{L}_{P^{(h)}}\subset\mathbb{C}^{-}$, we prove that the system
$\left[A_{(h)},C_{1,(h)},\dots,C_{r,(h)};B_{2},D_{1,2},\dots,D_{r,2}\right]$
@@
$\tilde{\varTheta}^{(h)}=\begin{bmatrix}\tilde{\varTheta}_{0}^{(h)}&\dots&\tilde{\varTheta}_{r}^{(h)}\end{bmatrix}\in\mathbb{R}^{n\times[m_{1}(r+1)]},$
@@
Then we can find the system $(A_{(h)}+\tilde{\varTheta}^{(h)}E_{0,(h)},C_{1,(h)}+\tilde{\varTheta}^{(h)}E_{1,(h)},\dots,C_{r,(h)}+\tilde{\varTheta}^{(h)}E_{r,(h)})$ associated with $\mathcal{L}^{(h-1,h)}$ is mean-square stable. This means the system $E_{0,(h)},E_{1,(h)},\dots,E_{r,(h)};A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)}]$ is stochastically detectable. Similarly, the Eq. (32) admits a unique stabilizing solution $Z^{(l)}$ such that $R_{22}(P^{(l)}+Z^{(l)})\succ 0$ and we also get $\text{Spec}\,\mathcal{L}^{(h,h+1)}\subset\mathbb{C}^{-}$ by using Proposition 3.7.
Since $\text{Spec}\,\mathcal{L}_{P^{(h)}}\subset\mathbb{C}^{-}$, from (i) and (ii) in Lemma 3.8, we get $\tilde{P}_{L}\succeq P^{(h)}+Z^{(h)}$ and $\text{Spec}\,\mathcal{L}_{P^{(h+1)}}\subset\mathbb{C}^{-}$. Similarly, by combining the conditions $R_{22}(P^{(h)}+Z^{(h)})\succ 0$ and $R(P^{(h)}+Z^{(h)})\preceq R(\tilde{P}_{L})$, we apply Corollary 4.5 from [16] to derive $R_{11}(P^{(h)}+Z^{(h)})\preceq\mathbb{R}_{22}^{\sharp}(P^{(h)}+Z^{(h)})\preceq\mathbb{R}_{22}^{\sharp}(\tilde{P}_{L})\prec 0$. Consequently, $P^{(h)}+Z^{(h)}\in\text{Dom }\mathcal{G}$. Substituting (32) into $\mathcal{G}\left(P^{(h)}+Z^{(h)}\right)$ by using Proposition 3.3, then gives:
$\mathcal{G}\left(P^{(h)}+Z^{(h)}\right)=(N_{1,(h)}-R_{12}(P^{(h)}+Z^{(h)})R_{22}(P^{(h)}+Z^{(h)})^{-1}N_{2,(h)})^{\top}R^{\sharp}_{22}(P^{(h)}+Z^{(h)})^{-1}$
$\quad\times(N_{1,(h)}-R_{12}(P^{(h)}+Z^{(h)})R_{22}(P^{(h)}+Z^{(h)})^{-1}N_{2,(h)}).$
@@
In this recursive process, the sequence $\{P^{(k)}\}_{k\geq 0}$ is monotonically non-decreasing and bounded above, so the sequence $\{P^{(k)}\}_{k\geq 0}$ is convergent and $\lim_{k\rightarrow\infty}Z^{(k)}=0$. Set $P^{*}=\lim_{k\rightarrow\infty}P^{(k)}$, given that
$\mathcal{G}(P^{*})=\mathcal{G}(\lim_{k\rightarrow\infty}P^{(k)})=\mathcal{G}\left(\sum_{s=0}^{\infty}Z^{(s)}\right)=\sum_{s=h+1}^{\infty}Z^{(s)}A_{(h)}+\sum_{s=h+1}^{\infty}A_{(h)}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}C_{l,(h)}^{\top}Z^{(s)}C_{l,(h)}$
$-\left(\sum_{s=h+1}^{\infty}B_{2}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(s)}C_{l,(h)}\right)^{\top}\left(R_{22}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(s)}D_{l,2}\right)^{-1}$
$\quad\times\left(\sum_{s=h+1}^{\infty}B_{2}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(s)}C_{l,(h)}\right)+\mathcal{G}\left(\sum_{s=0}^{h}Z^{(s)}\right),\forall h=0,1,2,\cdots.$
Since $\lim_{h\rightarrow\infty}\sum_{s=h+1}^{\infty}Z^{(s)}=0$ and $\lim_{h\rightarrow\infty}\mathcal{G}\left(\sum_{s=1}^{h}Z^{(s)}\right)=$
$\lim_{h\rightarrow\infty}(N_{1,(h)}-R_{12}(P^{(h+1)})R_{22}(P^{(h+1)})^{-1}N_{2,(h)})^{\top}R^{\sharp}_{22}(P^{(h+1)})^{-1}(N_{1,(h)}-R_{12}(P^{(h+1)})R_{22}(P^{(h+1)})^{-1}N_{2,(h)})$
@@
$0=P(A+B_{2}\tilde{L})+(A+B_{2}\tilde{L})^{\top}P+\sum_{l=1}^{r}(C+D_{l,2}\tilde{L})^{\top}P(C+D_{l,2}\tilde{L})+Q+\tilde{L}^{\top}R_{22}\tilde{L}+\tilde{L}^{\top}S_{2}+S_{2}^{\top}\tilde{L}-$ (33)
$[B^{\top}_{\tilde{W}}P+\sum_{l=1}^{r}D_{l\tilde{W}}^{\top}P(C+D_{l,2}\tilde{L})+S_{\tilde{L}\tilde{W}}]^{\top}(R_{\tilde{W}}+\sum_{l=1}^{r}D_{l\tilde{W}}^{\top}PD_{l\tilde{W}})^{-1}[B^{\top}_{\tilde{W}}P+\sum_{l=1}^{r}D_{l\tilde{W}}^{\top}P(C+D_{l,2}\tilde{L})+S_{\tilde{L}\tilde{W}}$],

### v2

@@ 3.3 Convergence Analysis of Iterative Sequence
Before presenting the next theorem, we first introduce two families of linear operators: $\mathcal{L}^{*}_{P^{(k)}}$ and $\mathcal{L}^{(k,k+1)}$ ($k=0,1,2,\ldots$), which are associated with the iteration sequences $\{Z^{(k)}\}_{k\geq 0}$ and $\{P^{(k)}\}_{k\geq 0}$ defined in (2.3).
$\mathcal{L}^{*}_{P^{(k)}}(Y)=\,Y(A_{(0)}+B_{1}\hat{K}_{1}(P^{(k)})+B_{2}L)+(A_{(0)}+B_{1}\hat{K}_{1}(P^{(k)})+B_{2}L)^{\top}Y+$ (24)
$\sum_{l=1}^{r}(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P^{(k)})+D_{l,2}L)^{\top}Y(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P^{(k)})+D_{l,2}L),\forall\,Y\in\mathbb{S}^{n},$
$\mathcal{L}^{(k,k+1)}(Y)=\,Y\left(A_{(k)}+B_{2}T_{(k+1)}\right)+\left(A_{(k)}+B_{2}T_{(k+1)}\right)^{\top}Y$ (25)
$+\sum_{l=1}^{r}\left(C_{l,(k)}+D_{l,2}T_{(k+1)}\right)^{\top}Y\left(C_{l,(k)}+D_{l,2}T_{(k+1)}\right),\forall\,Y\in\mathbb{S}^{n},$
where $T_{(k+1)}=-R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}$, $A_{(k)},C_{l,(k)}(1\leq l\leq r)$ and $N_{2,(k)}$ are defined in (9) and (10).
@@
  • There exist matrices $E_{l,(0)}(0\leq l\leq r)$ such that $\sum_{l=0}^{r}E_{l,(0)}^{\top}E_{l,(0)}=Q-S^{\top}(0)R(0)^{-1}S(0)$ and the system
    $\left[E_{0,(0)},E_{1,(0)},\dotsb,E_{r,(0)};A_{(0)},C_{1,(0)},\dotsb,C_{r,(0)}\right]$
@@
(i). The sequences $\{Z^{(k)}\}_{k\geq 0}$, $\{P^{(k)}\}_{k\geq 0}$ are well defined by (6) (7) (8) , and for each $k=0,1,2\ldots$ the following items are fulfilled:
@@
    $\operatorname{Spec}\,\mathcal{L}^{(k,k+1)}\subset\mathbb{C}^{-}$;
@@
    Let $L\in\mathcal{A}$ be arbitrary but fixed and $\tilde{P}_{L}$ be the stabilizing solution to the corresponding ARE (5), then $\tilde{P}_{L}\succeq P^{(k)}+Z^{(k)}$ and $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(k+1)}}\subset\mathbb{C}^{-}$;
@@
    $P^{(k)},P^{(k)}+Z^{(k)}\in\mathrm{Dom}\,\mathcal{G}$ and $\mathcal{G}(P^{(k)}+Z^{(k)})=$
    $-\left(N_{1,(k)}-R_{12}(P^{(k)}+Z^{(k)})R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}\right)^{\top}$
    $\times\mathbb{R}_{22}^{\sharp}(P^{(k)}+Z^{(k)})^{-1}$
    $\times\left(N_{1,(k)}-R_{12}(P^{(k)}+Z^{(k)})R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}\right),$
    where $\mathbb{R}_{22}^{\sharp}(P^{(k)}+Z^{(k)}),N_{1,(k)}$ and $N_{2,(k)}$ are defined in (10);
(ii). $\lim_{k\rightarrow\infty}P^{(k)}=\tilde{P}$, where $\tilde{P}$ is the stabilizing solution of the stochastic GTARE (3).
@@
Step 1. When $k=0$ , $P^{(0)}=0$ and $Z^{(0)}$ is the stabilizing solution to the following ARE:
$ZA_{(0)}+A_{(0)}^{\top}Z+\sum_{l=1}^{r}C_{l,(0)}^{\top}ZC_{l,(0)}+Q-S^{\top}(0)R(0)^{-1}S(0)-(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(0)})^{\top}$ (26)
$\quad\times(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,2})^{-1}(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(0)})=0.$
To show that $Z^{(0)}$ is well-defined, we first prove the existence of a unique stabilizing solution to the above ARE.
Given that the set $\mathcal{A}$ is nonempty, let $L\in\mathcal{A}$ be arbitrary but fixed and set $L^{(0)}=L$. Then we can find the system $(A_{(0)}+B_{2}L^{(0)},C_{1,(0)}+D_{1,2}L^{(0)},\dotsb,C_{r,(0)}+D_{r,2}L^{(0)})$ associated with the system $(A_{L},C_{1L},\dotsb,C_{rL})$ is mean-square stable. This means the system
@@
is mean-square stabilizable. Furthermore, the assumptions $R_{22}\succ 0$ and $\sum_{l=0}^{r}E^{\top}_{l,(0)}E_{l,(0)}=Q-S^{\top}(0)R(0)^{-1}S(0)\\ \succeq 0$ ensures that
@@
And the system $E_{0,(0)},E_{1,(0)},\,\,\dotsb\,\,,E_{r,(0)};A_{(0)},C_{1,(0)},\,\,\dotsb\,\,,C_{r,(0)}]$ is stochastically detectable. By virtue of Proposition 3.7, the ARE (26) admits a unique stabilizing solution $Z^{(0)}$ such that $R_{22}(Z^{(0)})\succ 0$. Thus, $Z^{(0)}$ is well-defined as the unique stabilizing solution to the ARE (26). This further implies that $\operatorname{Spec}\,\mathcal{L}^{(0,1)}\subset\mathbb{C}^{-}$.
Let $\tilde{P}_{L}$ denote the stabilizing solution to the ARE (5) associated with $L$. Since the operator $\mathcal{L}^{*}_{P^{(0)}}$ is associated with the system $(A_{L},C_{1L},\dotsb,C_{rL})$ is mean-square stable, we have the $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(0)}}\subset\mathbb{C}^{-}$. From (i) and (ii) in Lemma 3.8, we get $\tilde{P}_{L}\succeq P^{(0)}+Z^{(0)}$ and $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(1)}}\subset\mathbb{C}^{-}$.
Combining $R_{22}(\tilde{P}_{L})\succeq R_{22}\succ 0$ and $R_{11}\preceq R_{11}(\tilde{P}_{L})\prec 0$, we have $P^{(0)}$ and $\tilde{P}_{L}\in\mathrm{Dom}\,\mathcal{G}$. Since $R_{22}(P^{(0)}+Z^{(0)})\succeq R_{22}\succ 0$ and $R_{11}(P^{(0)}+Z^{(0)})\prec R_{11}(\tilde{P}_{L})\prec 0$, we obtain that $P^{(0)}+Z^{(0)}\in\mathrm{Dom}\,\mathcal{G}$.
Substituting the ARE (26) into $\mathcal{G}(P^{(0)}+Z^{(0)})$ by using Proposition 3.3, then gives $\mathcal{G}(P^{(0)}+Z^{(0)})=(N_{1,(0)}-R_{12}(Z^{(0)})R_{22}(Z^{(0)})^{-1}N_{2,(0)})^{\top}R^{\sharp}_{22}(Z^{(0)})^{-1}(N_{1,(0)}-R_{12}(Z^{(0)})R_{22}(Z^{(0)})^{-1}N_{2,(0)})$.
@@
Step 2. Assume k=h1k=h-1, $Z^{(h-1)}$ is well-defined and that $a_{h-1}-c_{h-1}$ hold. We shall prove that $Z^{(h)}$ is well-defined and that $a_{h}-c_{h}$ hold.
@@
$ZA_{(h)}+A_{(h)}^{\top}Z+\sum_{l=1}^{r}C_{l,(h)}^{\top}ZC_{l,(h)}+\sum_{l=0}^{r}E^{\top}_{l,(h)}E_{l,(h)}-(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(h)})^{\top}$ (27)
$\quad\times(R_{22}(P^{(h)})+\sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,2})^{-1}(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(h)})=0,$
where $E_{l,(h)}\in\mathbb{R}^{[m_{1}(r+1)]\times n}(0\leq l\leq r)$, and
$E_{l,(h)}=\begin{bmatrix}\mathbb{O}_{m_{1}\times n}^{\times(l)}\\ \sqrt{-\frac{1}{r+1}\mathbb{R}_{22}^{\sharp}\left(P^{(h)}\right)^{-1}}\left(N_{1,(h-1)}-R_{12}(P^{(h)})R_{22}(P^{(h)})^{-1}N_{2,(h-1)}\right)\\ \mathbb{O}_{m_{1}\times n}^{\times(r-l)}\end{bmatrix}$.
We demonstrate that $Z^{(h)}$ is well-defined. By the Proposition 3.7, proving that $Z^{(h)}$ is well-defined is equivalent to proving that the system
$\left[A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)};B_{2},D_{1,2},\dotsb,D_{r,2}\right]$
is mean-square stabilizable and the system
$\left[E_{0,(h)},E_{1,(h)},\dotsb,E_{r,(h)};A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)}\right]$
is stochastically detectable.
To prove the mean-square stabilizability of the system
$A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)};B_{2},D_{1,2},\dotsb,D_{r,2}],$
it suffices to set $L^{(h)}=L-\hat{K}_{2}(P^{(h)})$, where $\hat{K}_{2}(P)$ is defined in (16). We can then identify the system
$(A_{(h)}+B_{2}L^{(h)},C_{1,(h)}+D_{1,2}L^{(h)},\dotsb,C_{r,(h)}+D_{r,2}L^{(h)})$
associated with the operator $\mathcal{L}^{*}_{P^{(h)}}$. Given that the condition $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(h)}}\subset\mathbb{C}^{-}$, we thereby prove that the system
$\left[A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)};B_{2},D_{1,2},\dotsb,D_{r,2}\right]$
@@
$\tilde{\varTheta}^{(h)}=\begin{bmatrix}\tilde{\varTheta}_{0}^{(h)}&\dotsb&\tilde{\varTheta}_{r}^{(h)}\end{bmatrix}\in\mathbb{R}^{n\times[m_{1}(r+1)]},$
@@
Then we can find the system
$(A_{(h)}+\tilde{\varTheta}^{(h)}E_{0,(h)},C_{1,(h)}+\tilde{\varTheta}^{(h)}E_{1,(h)},\dotsb,C_{r,(h)}+\tilde{\varTheta}^{(h)}E_{r,(h)})$
associated with $\mathcal{L}^{(h-1,h)}$ is mean-square stable. This means the system
$E_{0,(h)},E_{1,(h)},\dotsb,E_{r,(h)};A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)}]$
is stochastically detectable. Similarly, from Proposition 3.7, the ARE (27) admits a unique stabilizing solution $Z^{(h)}$ such that $R_{22}(P^{(h)}+Z^{(h)})\succ 0$ and we also get $\operatorname{Spec}\,\mathcal{L}^{(h,h+1)}\subset\mathbb{C}^{-}$.
Since $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(h)}}\subset\mathbb{C}^{-}$, from (i) and (ii) in Lemma 3.8, we get $\tilde{P}_{L}\succeq P^{(h)}+Z^{(h)}$ and $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(h+1)}}\subset\mathbb{C}^{-}$. Similarly, combining $R_{22}(P^{(h)}+Z^{(h)})\succ 0$ and $R_{11}(P^{(h)}+Z^{(h)})\preceq R_{11}(\tilde{P}_{L})\prec 0$, we obtain that $P^{(h)}+Z^{(h)}\in\mathrm{Dom}\,\mathcal{G}$. Substituting (27) into $\mathcal{G}(P^{(h)}+Z^{(h)})$ by using Proposition 3.3, then gives $\mathcal{G}(P^{(h)}+Z^{(h)})=(N_{1,(h)}-R_{12}(P^{(h)}+Z^{(h)})R_{22}(P^{(h)}+Z^{(h)})^{-1}N_{2,(h)})^{\top}R^{\sharp}_{22}(P^{(h)}+Z^{(h)})^{-1}(N_{1,(h)}-R_{12}(P^{(h)}+Z^{(h)})R_{22}(P^{(h)}+Z^{(h)})^{-1}N_{2,(h)})$.
@@
Step 3. In this recursive process, the sequence $\{P^{(k)}\}_{k\geq 0}$ is monotonically non-decreasing and bounded above, so the sequence $\{P^{(k)}\}_{k\geq 0}$ is convergent and
$\lim_{h\rightarrow\infty}\sum_{s=h+1}^{\infty}Z^{(s)}=0,\lim_{k\rightarrow\infty}Z^{(k)}=0.$
Set $P^{*}=\lim_{k\rightarrow\infty}P^{(k)}$, given that
$\mathcal{G}(P^{*})=\mathcal{G}(\lim_{k\rightarrow\infty}P^{(k)})=\mathcal{G}\left(\sum_{s=0}^{\infty}Z^{(s)}\right)=$
$\sum_{s=h+1}^{\infty}Z^{(s)}A_{(h)}+\sum_{s=h+1}^{\infty}A_{(h)}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}C_{l,(h)}^{\top}Z^{(s)}C_{l,(h)}$
$-\left(\sum_{s=h+1}^{\infty}B_{2}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(s)}C_{l,(h)}\right)^{\top}\left(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}P^{*}D_{l,2}\right)^{-1}$
$\quad\times\left(\sum_{s=h+1}^{\infty}B_{2}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(s)}C_{l,(h)}\right)+\mathcal{G}\left(\sum_{s=0}^{h}Z^{(s)}\right),\forall h=0,1,2,\dotsb.$
Since $\lim_{h\rightarrow\infty}\sum_{s=h+1}^{\infty}Z^{(s)}=0$ and $\lim_{h\rightarrow\infty}\mathcal{G}\left(\sum_{s=0}^{h}Z^{(s)}\right)=$
$\lim_{h\rightarrow\infty}(N_{1,(h)}-R_{12}(P^{(h+1)})R_{22}(P^{(h+1)})^{-1}N_{2,(h)})^{\top}R^{\sharp}_{22}(P^{(h+1)})^{-1}$
$\times(N_{1,(h)}-R_{12}(P^{(h+1)})R_{22}(P^{(h+1)})^{-1}N_{2,(h)})=0,$
it follows that $\mathcal{G}(P^{*})=0$ and $P^{*}\succeq 0$ is the solution to the stochastic GTARE (3).
@@
By Proposition 3.3, expanding $\mathcal{G}(P^{*})$ via $P^{*}=(P^{*}-\tilde{P})+\tilde{P}$ yields: